### v3

# Enhancing the ODMR Signal of Organic Molecular Qubits

Abstract
In quantum information science and sensing, electron spins are often purified into a specific polarisation through an optical-spin interface, a process known as optically-detected magnetic resonance (ODMR). Diamond-NV centres and transition metals are both excellent platforms for these so-called colour centres, while metal-free molecular analogues are also gaining popularity for their extended polarisation lifetimes, milder environmental impacts, and reduced costs. In our earlier attempt at designing such organic high-spin $\pi$-diradicals, we proposed to spin-polarise by shelving triplet $M_{S}=\pm 1$ populations as singlets. This was recently verified by experiments albeit with low ODMR contrasts of $<1\%$ at temperatures above 5 K. In this work, we propose to improve the ODMR signal by moving singlet populations back into the triplet $M_{S}=0$ sublevel, designing a true carbon-based molecular analogue to the NV centre. Our proposal is based upon transition-orbital and group-theoretical analyses of beyond-nearest-neighbour spin-orbit couplings, which are further confirmed by ab initio calculations of a realistic trityl-based radical dimer. Microkinetic analyses point towards high ODMR contrasts of around $30\%$ under experimentally-feasible conditions, a stark improvement from previous works. Finally, in our quest towards ground-state optically-addressable molecular spin qubits, we exemplify how our symmetry-based design avoids Zeeman-induced singlet-triplet mixings, setting the scene for realising electron spin qubit gates.
Introduction
Polarised electron spins are promising quantum bit (qubit) candidates [1] with applications in quantum sensing [2] and quantum information science [3]. Purifying these spin magnetic dipoles into a particular polarisation requires irreversible decays and these have been achieved in solid-state spin defects [4, 5, 6, 7, 8, 9, 10, 11, 12, 13] by coupling the microwave-addressable electron spins with their orbital degrees of freedom [Fig. Introductiona]. Because the latter lies high in the optical regime, this technique has become known as optically-detected magnetic resonance (ODMR) – “detected” because the same optical-spin interface also enables readout of the spin polarisation. Platforms hosting optically addressable spins are then referred to as colour centres.
While early endeavours in this direction have focused on solid-state defects such as nitrogen-vacancy (NV) centres in diamond [14], their poor scalability and tunability have motivated the community to shift towards molecular platforms [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Transition metal complexes with high-spin ground states (GSs) offer the most natural starting point and this has been successful in Cr(IV), V(III), Ni(II), and, most recently, Ir(IV) complexes [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. However, in none of these systems were the spin lifetimes as long as NV centres, essential for applications like qubit operations, and this has been attributed to the larger spin-orbit couplings (SOCs) introduced by metals. Other challenges include higher costs and poorer sustainability. This has prompted a search for a molecular NV-centre mimic that is metal-free, just like diamond itself, notwithstanding the benefits of metal-based qubit systems such as higher biocompatibility [38, 39] and simpler protocols for multiqubit addressability [40, 41, 16, 42].
As a first step along this path, organic molecules were designed to demonstrate ODMR in high-spin excited states while acknowledging the limited lifetimes of these electronic excitations [43, 44, 45, 46, 47, 48, 49, 50, 51]. Subsequently, our group theoretically proposed a class of organic $\pi$-diradicals with optically-addressable high-spin ground states [52] that were also experimentally validated by Chowdhury et al. [53] and Kopp et al. [54]. In all of these systems, spin polarisation was likely attained by shelving the triplet $M_{S}=\pm 1$ population in the singlet manifold via spin-selective intersystem crossings (ISCs) while utilising the presence of additional singlet charge-recombined states absent from the triplet manifold [Fig. Introductionb]. (Interestingly, such imbalances have recently been proposed as evidence of low-energy electronic spin isomers [55].) Missing from the ODMR cycles are the ISCs between singlets and triplets selective for the $M_{S}=0$ triplet sublevel, which the NV centre exhibits. As mentioned in our previous work [52], this limits the maximum possible triplet $M_{S}=0$ population to $25\%$ at steady state, constraining the ODMR resolution. Moreover, this increases the risk of metastable singlet molecules returning as $M_{S}=\pm 1$ triplets, which would negate the accumulated spin polarisation. Indeed, in the aforementioned experimental works, poor ODMR contrasts of $<1\%$ were observed at temperatures above 5 K. We note in passing that, apart from $\pi$-diradicals, alkaline earth metal complexes [56, 57] and nitrenes [58] also display diradical properties and their functions as molecular colour centres have not been fully explored.
Our earlier theoretical model [52] had focused primarily on nearest-neighbour-only (through-bond-only) SOCs, which predicted ISCs to have $\Delta M_{S}=0$ (±$\pm 1$) selectivity for electron spins quantised parallelly (perpendicularly) to the dimer linkage. However, beyond-nearest-neighbour SOC effects have been shown to be experimentally relevant in ODMR, as in the recent work by Kopp et al. [54] [Fig. Introductionb], and such effects can induce ISCs with an opposite spin selectivity. Motivated by this observation, in this work, we design another class of organic $\pi$-diradicals, this time capable of ISCs involving both triplet $M_{S}=\pm 1$ and $M_{S}=0$ sublevels of different electronic states as facilitated by beyond-nearest-neighbour and nearest-neighbour SOCs respectively [Fig. Introductionc]. This way, we create a true analogue of the NV centre, which we predict to demonstrate enhanced ODMR signals by optically pumping the near-degenerate triplet and singlet local excitations. Our design principle stems from a careful group-theoretical analysis of the transitioning orbitals, which coincidentally also solves the problem of singlet-triplet mixing during electron paramagnetic resonance (EPR) measurements of weakly-coupled diradicals (as most $\pi$-diradicals are [59, 54]). As a realistic prototypical example of our design, the electronic structure of PT2TM-p-PT2TM [Fig. Introductionc] was shown at both density functional theory (DFT) and multi-configurational (MC) levels of theory to demonstrate an ODMR mechanism similar to the NV centre. The potential for signal enhancement through this ODMR pathway is supported by microkinetic analyses, where we found high ODMR contrasts of around $30\%$ to be experimentally possible (as compared to $<1\%$ in earlier theoretical [52] and experimental [53, 54] studies). In this regime, the steady-state triplet population exceeds $25\%$, which is the upper bound of prior works [52, 53, 54]. While not explored by the present work, we expect further mesitylation [Fig. Introductionc] to improve the luminescence quantum yield via geometrical relaxation [60, 61], as exemplified by Chowdhury et al. [53] to produce near-unity photoluminescence yields. This paves the way towards better designs of optically-addressable organic spins. Note that the alternacy symmetry of PT2TM-p-PT2TM is not important to our design, which should be separated from our earlier work [52].
FIG. 1: : (a) Simplified ODMR mechanism typical of diamond-NV defects. (b) Simplified ODMR mechanism of alternant $\pi$-diradicals as explored by our previous work [52]. Note that the CT-to-GS ISC rates are small if the two benzylic radicals are tethered at their meta positions, which is avoided by our new design (see main text). (c) ODMR mechanism of our new design, which more closely resembles the spin polarisation pathway of NV centres. Also drawn are the PT2TM-p-PT2TM diradical and its mesitylated counterpart, both of which satisfy our design requirements. (Abbreviations of electronic transitions: Abs = absorption; PL = photoluminescence; IC = internal conversion; ISC = intersystem crossing. Labels for electronic states: GS = ground state; LE = local excitation; CT = charge transfer. These labels are inherited from our previous analysis of the $\pi$-diradical electronic structure [52] with subscripts $A$, $B_{1}$, $B_{2}$, and $B_{3}$ signifying the state irreps.)
Results and Discussions
Investigating the role of beyond-nearest-neighbour SOCs
As we have previously shown [52], when two benzylic radicals are covalently connected with a significant torsion (as is the way most $\pi$-diradicals are experimentally constructed [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 59]), its SOC matrix elements are dominated by nearest-neighbour interactions between the two $2p$ atomic orbitals (AOs) of opposite $\pi$-rings (reason: El-Sayed rule  [^1] The El-Sayed rule states that the SOC matrix elements are the largest when the transitioning orbitals adopt different orientations [El-Sayed1963].) that are closest to the dimer linkage (reason: SOC is a local effect). Henceforth, we refer solely to $\pi$ AOs. Furthermore, if the spin quantisation axis lies parallel (perpendicular) to the line connecting the two $2p$ AO centres, then the ISC spin selectivity arising from the SOC matrix elements is strictly $\Delta M_{S}=0$ (±$\pm 1$) due to symmetry [52]. Note that the last condition is satisfied for interactions at the dimer bond because the spin axes (wihout applying any magnetic fields) coincide with the dimer’s high symmetry axes. However, ISCs in $\pi$-diradicals occur not between pairs of $2p$ AOs but between pairs of delocalised molecular orbitals (MOs) that are each a linear combination of a few $2p$ AOs. When one of the transitioning MOs contains a node at the dimer linkage, the relevant SOC matrix elements necessarily involve next-nearest-neighbour couplings not considered by the above analysis. Such effects are generally small following the local nature of SOCs, yet they can be appreciable if most of the transitioning MO’s amplitude is piled up at these next-nearest-neighbour sites (as opposed to nearest-neighbour couplings weighted by small MO amplitudes). Importantly, next-nearest-neighbour interactions involve two $2p$ AO centres that are not colinear with the spin quantisation axis, causing the ISC selectivity to break down [Fig. Investigating the role of beyond-nearest-neighbour SOCsa]. Indeed, for two benzylic radicals covalently tethered at the para positions with $90^{\text{o}}$ torsion (hereafter called benzyl-p-benzyl), the next-nearest-neighbour SOC evaluated over singly-occupied $2p$ AOs biases the $\Delta M_{S}=\pm 1$ transitions with $68.1\%$ selectivity (noting that the spin quantisation ($z$-)axis lies parallel to the dimer linkage due to the molecular $D_{2d}$ point group symmetry) and this SOC has a significant amplitude of $0.13\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. For comparsion, the nearest-neighbour analogue has a perfect spin selectivity of $\Delta M_{S}=0$ with a larger amplitude of $1.37\mathrm{\,cm^{-1}}$. Note that these SOC matrix elements represent upper bounds to the true molecular values because we have assumed one full electron in each AO.
FIG. 2: : (a) Breakdown of the spin selection rule for ISCs involving next-nearest-neighbour SOCs. (b) Amplitudes of the SOC matrix elements evaluated between two carbon $2p$ AOs centred on opposite $\pi$-systems of the benzyl-p-benzyl diradical. Each $2p$ AO is assumed to be occupied by one full electron with either the same or opposite spin as its counterpart on the other benzylic fragment. Taking the difference between the matrix elements-squared of the two spin alignments and weighing the result by their sums give the spin selectivity. (c) Sketches of the frontier Hückel MOs for the benzylic monoradical. Also illustrated are their parities under $C_{2}\left(z\right)$ rotation. (d) Approximate electronic structure of benzyl-p-benzyl obtained from the Hückel theory analysis in (c). The excitation labels are inherited from Ref. [52] and the subscripts indicate the state irreps. In this work, all molecular geometries are presented in the coordinate frame depicted in (e).
Perfecting the spin selectivity using the molecular point group symmetry
That ISC in $\pi$-diradicals can actually occur with both spin selectivities is contrary to most theoretical analyses [80, 81, 82] (including ours [52]) and opens up the possibility of transferring the full electronic structure of NV centres into molecules. Our next step is to perfect the next-nearest-neighbour spin selectivity via symmetry considerations. We note that by connecting two benzylic radicals at their para positions, the resulting diradical has at least a $D_{2}$ point group symmetry with irreducible representations (irreps) organised by their characters under $C_{2}\left(y\right)$ and $C_{2}\left(z\right)$ rotations. The irrep of any electronic state is in turn the direct product of its orbital and spin irreps. Focusing on the orbital part, if we continue to align the $z$-axis along the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsb,e], then a $C_{2}\left(y\right)$ rotation becomes equivalent to swapping the two monomers. Therefore, for two weakly-interacting $\pi$-systems of significant torsion, the low-lying excitations are expected to appear in near-degenerate pairs of $\left\{A,B_{1}\right\}$ and $\left\{B_{2},B_{3}\right\}$, each comprising a state of opposite parity under $C_{2}\left(y\right)$, i.e. equal probability of exciting either monomer. These two pairs can be distinguished by their transformations under $C_{2}\left(z\right)$ and, for that, it is easier to think in the basis of monomeric $\pi$ MOs. We first note that $2p$ AOs have $\pi$ symmetry, that is, their phases change under $C_{2}$ rotation about the AO centre. Further, for the two $2p$ AOs at the dimer linkage, that $\pi$-symmetry-defining $C_{2}$ rotation coincides with the diradical’s $C_{2}\left(z\right)$ symmetry element. Hence, a monomeric MO with a non-zero amplitude at the dimer linkage must be antisymmetric under $C_{2}\left(z\right)$ rotation. By contrast, for a monomeric MO to be symmetric under $C_{2}\left(z\right)$ rotation, it must possess a node at the dimer linkage. For instance, the Hückel method predicts the highest occupied, singly occupied and lowest unoccupied MOs (abbreviated as HOMO, SOMO, and LUMO respectively) of the benzylic radical to have $+1$, $-1$, and $+1$ characters under $C_{2}\left(z\right)$ rotation [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Finally, we recall that the symmetry of the orbital wavefunction is the direct product of the symmetries of all singly occupied orbitals. Therefore, the ground states of benzyl-p-benzyl, having two unpaired electrons in the SOMOs, transform as either $A$ or $B_{1}$. Similarly, the lowest charge transfer (CT) singlets with all doubly-occupied orbitals (also known as charge-recombined or zwitterionic states of the two radicals) also transform as either $A$ or $B_{1}$. As for the lowest local excitations (LEs), they are a linear combination of HOMO-to-SOMO and SOMO-to-LUMO excitations (an outcome of alternancy symmetry; see Ref. [52]) and hence transform as either $B_{2}$ or $B_{3}$. Including the characters under $C_{2}\left(y\right)$ rotation yields the approximate energy level diagram illustrated in Fig. Investigating the role of beyond-nearest-neighbour SOCsd (for a review of the electronic structure of benzylic diradicals, see our earlier work [52]).
Returning to our goal of attaining perfect ISC spin selectivities, we note that the singlet, triplet $M_{S}=0$, and triplet $M_{S}=\pm 1$ spin sublevels transform as $A$, $B_{1}$, and $\left\{B_{2},B_{3}\right\}$ respectively [83, 84]. For an ISC to be symmetry-allowed, the SOC matrix elements must contain the totally symmetric irrep. Hence, because the SOC Hamiltonian is totally symmetric, ISC processes in benzyl-p-benzyl are spin-selective for $\Delta M_{S}=\pm 1$ when moving between triplet LEs and singlet CTs, and $\Delta M_{S}=0$ when moving between singlet CTs and triplet GSs. Actually, we already knew this from the previous section: The first ISC occurs between SOMOs and LUMOs (or HOMOs) of opposite monomers and only the latter has a node at the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Thus, this process occurs predominantly via next-nearest-neighbour SOCs, which we found before to exhibit $\Delta M_{S}=\pm 1$ selectivity [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. Similarly, the second ISC connects between two SOMOs of opposite monomers and is mediated mostly by nearest-neighbour SOCs with $\Delta M_{S}=0$ selectivity [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. Yet, the key improvement achieved by this section is the perfection of ISC spin selectivities to $100\%$ by making the opposite spin channel symmetry-forbidden. From a microscopic perspective, we have destructively interfered any deleterious ISC channels using other higher-nearest-neighbour SOCs, treated more generally by the group-theoretical approach. Importantly, this electronic structure produces the ODMR mechanism shown in Fig. Introductionc, where populations in the triplet $M_{S}=\pm 1$ LEs are moved into the triplet $M_{S}=0$ GS via two sequential ISCs of opposite spin selectivities. In other words, such molecules exhibit the ODMR mechanism of NV centres.
As an aside, while it is tempting to consider only symmetry arguments and ignore the discussions made in the previous section, note that symmetry-allowed transitions need not occur with high amplitudes. Therefore, one still needs to check if the allowed ISCs are facilitated by either nearest- or next-nearest-neighbour SOCs, which have the largest amplitudes. For example, the meta-connected diradicals explored in our previous work [52] have negligible SOCs between singlet CTs and triplet GSs [Fig. Introductionb] because they are dominated by third-nearest-neighbour effects with a maximum SOC amplitude of $0.02\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. In other words, our plan for enhanced ODMR contrasts requires para-connectivity between the two benzylic radicals, consistent with our group-theoretical analysis (meta-connectivity does not yield $D_{2}$ point group symmetry). We note in passing that vibronic perturbations to the SOC operator, which are not considered by this study, can modify the spin selectivity as well [13, 85]. This will be explored in our upcoming studies.
Our proposal
We now introduce a molecular design that realises the ODMR mechanism described in Fig Introductionc:
  1. 1.
    The $\pi$-diradical should be a dimer centralised around a benzyl-p-benzyl-like framework. This ensures that the diradical’s electronic structure resembles Fig. Investigating the role of beyond-nearest-neighbour SOCsd.
  2. 2.
    Sufficient steric hindrances should be installed around the dimer linkage to maintain a $90^{\text{o}}$ torsion. Otherwise, para connectivity facilitates $\pi$-bonding and encourages a singlet ground state [86, 87, 88, 66, 89]. For instance, chloro substituents can be introduced at the meta positions, which is common practice when synthesising luminescent benzyl-type radicals [66].
  3. 3.
    Any additional $\pi$-conjugation to the benzylic radical should continue to localise the lowest-lying excitations around the benzylic group. For instance, phenyl substituents added to the benzylic radical (typically used to improve radical stability and luminescence) should have less-extensive internal $\pi$-conjugation so that their lowest transitions are energetically separated from the benzylic fragment’s.
  4. 4.
    These extra substituents should also maintain the monoradical’s $C_{2}\left(z\right)$ symmetry, where the $z$-axis lies along the ipso-para direction. The reason is that the $C_{2}\left(z\right)$ axis is the only symmetry element discriminating between the monomeric $\pi$ MOs [Fig. Investigating the role of beyond-nearest-neighbour SOCsc], thus their parities under $C_{2}\left(z\right)$ dictate the spin selectivity of their intermonomer ISCs. (For the same reason, diradicals of $D_{2}$ point group symmetry are not the only molecules satisfying our design criteria. Non-dimeric diradicals can also be constructed with $C_{2}$ symmetry and the above near-degenerate pairs of $\left\{A,B_{1}\right\}$ and $\left\{B_{2},B_{3}\right\}$ excitations will become split in energy, thereby offering a ladder of states through which the singlet molecules can decay. Here, we will focus solely on dimeric diradicals of $D_{2}$ symmetry for simplicity.) As an example, when substituting the two hydrogens on the benzylic centre, both hydrogens should be replaced simultaneously and with identical moieties.
These features are summarised in Fig. Our proposala using our prototype PT2TM-p-PT2TM diradical as an example. Note that its alternacy symmetry is not crucial to our design, which should be distinguished from our earlier work [52].
FIG. 3: : (a) Features of our diradical design, demonstrated on the PT2TM-p-PT2TM diradical. (b) Energy level diagram of PT2TM-p-PT2TM, computed using TDDFT. Also included are the root-mean-squared SOC matrix elements, averaged over all channels. These calculations were performed using MCSCF/CI methods on the central tetra-chlorinated benzyl-p-benzyl fragment. (c) Ground state spin density (isovalue = 0.004) and natural transition orbitals (isovalue = 0.040) for the first few excited states of PT2TM-p-PT2TM, obtained using BS-DFT/TDDFT.
Ab initio calculations of the prototype PT2TM-p-PT2TM diradical
[…]
As speculated, the lowest LE transitions are largely localised on the central (tetra-chlorinated) benzyl-p-benzyl fragment [Fig. Our proposalc]. Hence, for computational efficiency, the SOC matrix elements were computed for this central fragment instead, capping any open ends with hydrogen. Indeed, the MCSCF/CI results suggest appreciable SOC matrix elements along the proposed ODMR pathway [Fig. Our proposalb]. While, strictly speaking, these calculations present only an upper bound to the true SOC matrix elements of the full PT2TM-p-PT2TM diradical, their values should not decrease by much upon weak $\pi$-conjugation to the additional trichlorophenyl substituents. Most importantly, these calculations lend credence to the theoretical proposal presented by this work.
Symmetry avoids Zeeman-induced singlet-triplet mixings during EPR spectroscopy
Interestingly, in connecting two identical monoradicals of (at least) $C_{2}$ point group symmetry to form a diradical of (at least) $D_{2}$ point group symmetry, the resulting structure can avoid undesirable singlet-triplet mixings found in many weakly-coupled diradicals under an applied magnetic field. This would alleviate most magnetic-field-induced decoherence events in the ground state [91, 92] (notice that the singlet and triplet GSs differ by either the coherence or a single spin flip [52]). It also facilitates Rabi nutation experiments in EPR spectroscopy [54], amongst other benefits.
To introduce the problem, we consider the spin Hamiltonian $\hat{H}_{\text{spin}}$ of two electron spins $\hat{{\bf S}}_{j}$ ($j=1,2$) under an applied magnetic field ${\bf B}$:
$\hat{H}_{\text{spin}}$ $=\mu_{\text{B}}{\bf B}\cdot{\bf g}_{1}\cdot\hat{{\bf S}}_{1}+\mu_{\text{B}}{\bf B}\cdot{\bf g}_{2}\cdot\hat{{\bf S}}_{2}+2J\hat{{\bf S}}_{1}\cdot\hat{{\bf S}}_{2}$
$\quad+\frac{\mu_{0}g_{\text{e}}^{2}\mu_{\text{B}}^{2}}{4\pi}\left[\frac{\hat{{\bf S}}_{1}\cdot\hat{{\bf S}}_{2}}{\left|{\bf r}\right|^{3}}-\frac{3\left(\hat{{\bf S}}_{1}\cdot{\bf r}\right)\left(\hat{{\bf S}}_{2}\cdot{\bf r}\right)}{\left|{\bf r}\right|^{5}}\right].$ (1)
The first two terms represent each spin’s interaction with the magnetic field (i.e. the spin Zeeman terms), characterised by different $g$-tensors ${\bf g}_{j}$ ($j=1,2$). The third term symbolises the exchange interaction between the two spins with coupling $2J$ and is responsible for singlet-triplet splittings. The last term denotes the spin-spin dipolar interaction that gives rise to zero-field splittings in organic diradicals; here, the spatial degrees of freedom (characterised by ${\bf r}$) have been integrated over some orbital subspace (for instance, the ground state). Finally, the symbols $\mu_{\text{B}}$, $\mu_{0}$, and $g_{\text{e}}$ represent, respectively, the Bohr magneton, the vacuum magnetic permeability, and the electron $g$-factor. Implicit to the above expression is the definition of a four-dimensional Hilbert space spanned by tensor products of the two electron spin states. By applying the angular momentum sum rules, one may also express the Hilbert space in the basis of one singlet and three triplet spin states [93, 91]. This amounts to re-expressing Eq. (1) in terms of the total spin operator $\hat{{\bf S}}\equiv\hat{{\bf S}}_{1}+\hat{{\bf S}}_{2}$, which yields
$\hat{H}_{\text{spin}}$ $=\mu_{\text{B}}{\bf B}\cdot{\bf g}_{1}\cdot\hat{{\bf S}}+\mu_{\text{B}}{\bf B}\cdot\left({\bf g}_{2}-{\bf g}_{1}\right)\cdot\hat{{\bf S}}_{2}$
$\quad+J\left(\hat{{\bf S}}^{2}-\hat{{\bf S}}_{1}^{2}-\hat{{\bf S}}_{2}^{2}\right)+\hat{{\bf S}}\cdot{\bf D}\cdot\hat{{\bf S}}.$ (2)
The procedure for transforming the last term of Eq. (1) into the last term of Eq. (2) is presented in most EPR texts [91, 92] with ${\bf D}$ being referred to as the zero-field splitting tensor. We now see the problem: If ${\bf g}_{1}\neq{\bf g}_{2}$, true in most organic systems, then the singlet-triplet basis block-diagonalises every term in Eq. (2) (with the triplets forming their own block) except for the second one. In other words, the second term induces mixings between the singlet state and the triplet manifold. This, as mentioned, leads to deleterious effects such as spin decoherence [91, 92] and additional Rabi frequencies [54]. One solution would be to engineer an exchange coupling $J$ that is large enough for the Zeeman-induced singlet-triplet couplings of $\sim\mu_{\text{B}}B\left|g_{2}-g_{1}\right|$ to be a perturbation relative to the singlet-triplet gap. However, as explained in our earlier work [52], further enhancement of the exchange coupling in $\pi$-diradicals is likely to result in a singlet GS, i.e. $J$ is typically antiferromagnetic [86, 87, 88, 66].
The alternative approach would be to somehow make elements that differ between ${\bf g}_{1}$ and ${\bf g}_{2}$ disappear under experimental conditions. This can be achieved in our diradicals by aligning the magnetic field ${\bf B}$ parallelly to the dimer linkage. (We acknowledge that aligning the molecules, say, in a liquid crystal, may be an experimental challenge, although this has been previously demonstrated [94, 95].) To see that, we first note that elements of the $g$-tensor have the following approximate expression, obtained from second-order perturbation theory in the SOC: [91, 92]
$g_{ab}$ $=g_{\text{e}}\delta_{ab}+2\lambda\sum_{\Psi^{\prime}}\frac{\langle\Psi|\hat{L}_{a}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|\hat{L}_{b}|\Psi\rangle}{E_{\Psi}-E_{\Psi^{\prime}}}.$ (3)
Here, $\lambda$ is an effective SOC constant and $\left\{|\Psi^{\prime}\rangle\right\}$, $\left\{E_{\Psi^{\prime}}\right\}$ respectively denote the orbital eigenfunctions and eigenenergies of the molecule’s non-relativistic electronic Hamiltonian, with $\Psi$ being the orbital eigenfunction under consideration (say, the ground state wavefunction). The indices $a$ and $b$ run over all three-dimensional coordinates (such as $x$, $y$, and $z$) and $\hat{L}_{a}$ is the $a$-th component of the orbital angular momentum operator $\hat{{\bf L}}$. Because each monoradical of our proposed system is symmetric about $C_{2}\left(z\right)$ rotation (using the coordinate frame defined by Fig. Investigating the role of beyond-nearest-neighbour SOCse, i.e. $z$-axis along dimer linkage), the $g$-tensor must remain invariant under $C_{2}\left(z\right)$ rotation. Such a rotation takes $\left\{\hat{L}_{x},\hat{L}_{y},\hat{L}_{z}\right\}\rightarrow\left\{-\hat{L}_{x},-\hat{L}_{y},\hat{L}_{z}\right\}$, hence, using Eq. (3), we obtain
${\bf g}$ $\equiv\left(\begin{array}[]{ccc}g_{xx}&g_{xy}&g_{xz}\\ g_{yx}&g_{yy}&g_{yz}\\ g_{zx}&g_{zy}&g_{zz}\end{array}\right)=\left(\begin{array}[]{ccc}g_{xx}&g_{xy}&-g_{xz}\\ g_{yx}&g_{yy}&-g_{yz}\\ -g_{zx}&-g_{zy}&g_{zz}\end{array}\right).$ (10)
By contradiction, it must be that
${\bf g}$ $=\left(\begin{array}[]{ccc}g_{xx}&g_{xy}&0\\ g_{yx}&g_{yy}&0\\ 0&0&g_{zz}\end{array}\right).$ (14)
Next, the two monoradicals in the dimer are related by a $C_{2}\left(y\right)$ rotation. Taking Eq. (14) to be the $g$-tensor of radical 1, we find the $g$-tensors of both radicals to be
${\bf g}_{1}$ $=\left(\begin{array}[]{ccc}g_{xx}&g_{xy}&0\\ g_{yx}&g_{yy}&0\\ 0&0&g_{zz}\end{array}\right),$ (18)
${\bf g}_{2}$ $=\left(\begin{array}[]{ccc}g_{xx}&-g_{xy}&0\\ -g_{yx}&g_{yy}&0\\ 0&0&g_{zz}\end{array}\right).$ (22)
Therefore, because both ${\bf g}_{1}$ and ${\bf g}_{2}$ have the same $g_{zz}$ components, the second term of Eq. (2) vanishes when the magnetic field ${\bf B}$ falls along the $z$-axis (the dimer linkage), i.e.
$\hat{H}_{\text{spin}}$ $=\mu_{\text{B}}g_{zz}\left|{\bf B}\right|\hat{S}_{z}$
$\quad+J\left(\hat{{\bf S}}^{2}-\hat{{\bf S}}_{1}^{2}-\hat{{\bf S}}_{2}^{2}\right)+\hat{{\bf S}}\cdot{\bf D}\cdot\hat{{\bf S}}.$ (23)
The singlet-triplet states may now block-diagonalise $\hat{H}_{\text{spin}}$ with no singlet-triplet mixings present. Furthermore, due to the approximate $D_{2d}$ point group symmetry of our system, the zero-field splitting caused by ${\bf D}$ also occurs along the $z$-axis. This may be further rationalised by the cylindrical spin density distribution of the diradical [92]. Therefore, the singlet-triplet basis fully-diagonalises $\hat{H}_{\text{spin}}$ in Eq. (23) when the magnetic field is applied along the dimer linkage.
Predicted ODMR contrasts
As a final theoretical piece to this puzzle, we construct a microkinetic model of the excited state kinetics and predict the ODMR contrast at steady state. For simplicity, we shall take the incoherent limit and consider only changes to the eigenstate populations. Also, populations of degenerate eigenstates are pooled into a single variable. Therefore, the relevant eigenstates are
  1. a.
    the singlet GS (labelled GS, degeneracy = 1),
  2. b.
    the triplet $M_{S}=0$ GS (labelled GT0 for ground triplet, degeneracy = 1),
  3. c.
    the triplet $M_{S}=\pm 1$ GS (labelled GT1, degeneracy = 2),
  4. d.
    the singlet CTs (labelled CT, degeneracy = 2),
  5. e.
    the singlet LEs (labelled ES for excited singlet, degeneracy = 2),
  6. f.
    the triplet $M_{S}=0$ LEs (labelled ET0 for excited triplet, degeneracy = 2), and
  7. g.
    the triplet $M_{S}=\pm 1$ LEs (labelled ET1, degeneracy = 4).
Here, we continue to assume a spin quantisation axis parallel to the dimer linkage. The kinetic processes connecting these eigenstates and their associated rate constants are presented in Fig. Predicted ODMR contrasts; there, the symbols adopt their usual meanings. As far as possible, model parameters were selected within experimental reason. Listed below are our justifications for these parameters and the reader may safely skip it at the first pass.
  1. 1.
    Also modelled are the vibrationally-excited LE states, labelled with an asterisk (*). These states represent the wavefunction following an initial Franck-Condon photoexcitation and will, within picoseconds, undergo vibrational relaxation to the ground vibrational level of the respective LE states, labelled without an asterisk ( ).
  2. 2.
    The experimentally-observed IC rates from LE to GS triplets are between $10^{8}$ and $10^{10}\mathrm{\,s^{-1}}$ [68, 75, 53, 54]. In our model, we shall pick the upper bound as a conservative estimate. Then, because experiments predict the IC rate from CT to GS singlets to be an order of magnitude slower [53, 54], we place its value at $10^{9}\mathrm{\,s^{-1}}$. Finally, IC from LE to CT singlets must be faster in view of the smaller energy gap [90] and we estimate this rate to be $10^{11}\mathrm{\,s^{-1}}$.
  3. 3.
[…]
  4. 4.
    The weakly-coupled pair of doublet spins in the GS can mix and decohere which, in the singlet-triplet basis, translates into population transfers among the singlet and triplet GSs. Therefore, we shall assume a uniform rate constant $k_{\text{decoh}}$ for mixing between the four near-degenerate magnetic levels (one from the singlet and three from the triplet). This is consistent with Kopp et al. [54], where the singlet to triplet GS ISCs occur at around the same timescale as the triplet spin-lattice and spin-spin relaxations. The same decoherence pathway is also expected among the singlet and triplet LEs. Special attention is given to mixing among the triplet spin sublevels, denoted by a rate constant $\kappa_{\text{decoh}}$. In our simulations, we will always set $\kappa_{\text{decoh}}=k_{\text{decoh}}$ unless a microwave drive is applied (necessary for finding the ODMR contrast). This introduces additional triplet spin mixing that, assuming a saturating microwave field, is mathematically equivalent to the limit of $\kappa_{\text{decoh}}\rightarrow\infty$, i.e. full mixing between the triplet sublevels [26].
Solving the kinetic model for the steady-state populations yields the ODMR contrast and optically-induced spin polarisation via the following expressions [11, 12, 13]:
ODMR contrast
$\quad=\frac{\sum_{j=0,1}n_{\text{ET}j}^{\text{ss},\infty}-\sum_{j=0,1}n_{\text{ET}j}^{\text{ss}}}{\sum_{j=0,1}n_{\text{ET}j}^{\text{ss}}}\times 100\%$ (24)
and
Spin polarisation $=\frac{n_{\text{GT0}}^{\text{ss}}-n_{\text{GT1}}^{\text{ss}}/2}{n_{\text{GT0}}^{\text{ss}}+n_{\text{GT1}}^{\text{ss}}/2}\times 100\%,$ (25)
where the factor of $1/2$ arises from the double degeneracy of the triplet $M_{S}=\pm 1$ GSs. Here, $n_{\Psi}^{\text{ss}}$ denotes the steady-state population of state $\Psi$ at $\kappa_{\text{decoh}}=k_{\text{decoh}}$, while $n_{\Psi}^{\text{ss},\infty}$ denotes the same quantity at $\kappa_{\text{decoh}}\rightarrow\infty$, i.e. under a microwave resonance. (The rate equations are available in Supplementary Information 1.)
[…]
FIG. 4: : Schematic diagram depicting the excited-state transitions being considered by our microkinetic model.
FIG. 5: : Steady-state solutions to the (a) optically-induced spin polarisation, (b) ODMR contrast, (c) triplet $M_{S}=0$ GS population, and (d) triplet $M_{S}=0$ LE population (which simulates the emission intensity). Plots are made across varying optical pump rates $k_{\text{abs}}$ and spin decoherence rates $k_{\text{decoh}}$ within experimental ranges. Other parameters are defined in Fig. Predicted ODMR contrasts, chosen to be as realistic as possible.
Conclusion
By covalently tethering two benzylic radicals at their para positions, the resulting diradical of $D_{2}$ point group symmetry has a triplet GS of irrep $B_{1}$, lowest singlet CTs of irreps $A$ and $B_{1}$, and lowest triplet LEs of irreps $B_{2}$ and $B_{3}$. Thus, by group-theoretical considerations, spin polarisation can be achieved via a $\Delta M_{S}=\pm 1$ ISC from LEs to CTs, followed by a $\Delta M_{S}=0$ ISC from CTs to GSs, just like in an NV centre. Moreover, the associated SOC matrix elements are appreciable due to next-nearest-neighbour effects for the former and nearest-neighbour effects for the latter. We expect this success to be generalisable to any diradical based upon two para-connected benzylic radicals. This is illustrated in the PT2TM-p-PT2TM diradical: Firstly, the benzylic radicals are meta-chlorinated to lock the inter-radical torsion by steric repulsion, ensuring that the GS comprises both triplets and singlets at near-degeneracy. Next, the stabilising phenyl rings are specially chosen to not break the $C_{2}$ rotational symmetry down the benzylic ipso-para direction, which is the essential symmetry element of the aforementioned benzylic para-dimer. Finally, every benzylic radical has one more chloro substituent than each of its phenyl branches, thereby localising the lowest-lying excitations on the tetra-chlorinated benzylic fragment. Indeed, these expectations are reflected in the ab initio results by producing the desired electronic structure for ODMR.
To our best knowledge, our new design has checked most boxes for a robust optical-spin interface [17]. Perhaps a minor problem not addressed by this work is the poor emissive properties of our PT2TM-p-PT2TM prototype due to its alternacy symmetry [98, 99, 52]. For that, we expect techniques of alternacy symmetry breaking [98, 99, 45] and excited-state symmetry breaking [60, 61] employed by the organic light-emitting diode (OLED) community to be useful. In fact, the latter was recently demonstrated by Chowdhury et al. [53], who synthesised trityl-based diradicals of near-unity luminescence quantum yields by mesitylating at the para positions. When applied to PT2TM-p-PT2TM, a possible diradical would look like Fig. Introductionc, which is a suitable synthetic starting point. Notably, the ODMR contrast for this novel spin polarisation pathway is expected to be around $30\%$ with more than $25\%$ of the population being in the triplet manifold at steady state, an unequivocal improvement from earlier theoretical [52] and experimental [53, 54] works. Regarding qubit operations, our design also overcomes problems of Zeeman-induced spin decoherence and observations of multiple Rabi frequencies during EPR measurements of weakly-coupled diradicals. These are all crucial steps towards the realisation of optically-addressable molecular spin qubits.
Methods
The geometry of benzyl-p-benzyl was relaxed as a triplet on UB3LYP-D3BJ/def2-SVP [100, 101] while enforcing $D_{2d}$ point group symmetry. Due to the lack of steric hindrances, the optimisation converged to a saddle point on the potential energy surface with a single imaginary frequency of amplitude $57.8\mathrm{\,cm^{-1}}$. These calculations were done using the ORCA 5.0 code [102]. We then computed the SOC matrix elements between carbon $2p$ AOs of opposite $\pi$-systems using the single-electron form of the SOC operator [103], taking each AO to be singly-occupied and using a carbon effective nuclear charge of 3.9 [104]. These integrals were computed using PySCF 2.5.0 [105] and the carbon $2p$ AOs were represented by the STO-6G basis set.
Hereafter, all ab initio calculations were performed with the ORCA 5.0 code [102]. Geometries of PT2TM-p-PT2TM and its central tetra-chlorinated fragment (labelled Cl4M-p-Cl4M) were optimised as triplets on UB3LYP-D3BJ/def2-SVP [100, 101] with harmonic vibrational frequency analyses done to ensure no imaginary frequencies. The resulting geometries had at least $D_{2}$ point group symmetries. For the full PT2TM-p-PT2TM diradical, its electronic structure was estimated by spin-unrestricted TDDFT/TDA at the UB3LYP/def2-SVPD level using both triplet and BS singlet ground states as reference. In cases where the latter approach yielded BS solutions of $\langle{\bf S}^{2}\rangle\approx 1$, we assumed negligible discrepancies to the singlet energies from spin contamination because the results were similar to the triplet excitations (within 0.1 eV), which are the most-probable spin contaminants [106]. Excited state irreps were then inferred from the polarisations of the respective transition dipoles. In performing a BS-DFT calculation, we had also obtained the open-shell singlet GS energy via the following expression from Yamaguchi et al. [106]:
$E_{\text{S}}-E_{\text{T}}$ $=\frac{\langle{\bf S}^{2}\rangle_{\text{T}}}{\langle{\bf S}^{2}\rangle_{\text{T}}-\langle{\bf S}^{2}\rangle_{\text{BS}}}\left(E_{\text{BS}}-E_{\text{T}}\right),$ (26)
where $E_{\text{T}}$, $E_{\text{BS}}$, and $E_{\text{S}}$ are the energies of the triplet, BS singlet, and open-shell singlet states while $\langle{\bf S}^{2}\rangle_{\text{T}}$ and $\langle{\bf S}^{2}\rangle_{\text{BS}}$ are the respective expectations of the total spin-squared. The smaller Cl4M-p-Cl4M fragment was then used to compute the SOC matrix elements at the MCSCF/CI level, also using the def2-SVPD basis set. For that, we first performed a complete active space self-consistent field (CASSCF) calculation with 10 active electrons and 10 active orbitals, state-averaging the energy over the ground triplet and singlet states (i.e. SA2). Thereafter, the excited states were estimated using the complete active space configuration interaction (CASCI) method, among which the SOC matrix elements were computed. In this work, the electron spin states are expressed in the symmetry-aligned coordinate frame shown in Fig. Investigating the role of beyond-nearest-neighbour SOCse. Finally, with the monoradical of PT2TM-p-PT2TM (labelled PT2TM-H), its geometry was first optimised as a doublet on UB3LYP-D3BJ/def2-SVP [100, 101] and the presence of a local minimum of $C_{2}$ point group symmetry was confirmed by harmonic vibrational frequency analysis. Then, using the def2-SVPD basis set, orbitals were optimised by CASSCF(11,11) state-specific to the ground doublet, following which the excited doublet energies were obtained by CASCI. In this last calculation, the RIJCOSX approximation (RIJCOSX = resolution of identity approximation for the Coulomb term and chain-of-spheres approximation for the exchange term) was applied using the def2/JK auxiliary basis set. All CASCI energies were corrected by the strongly contracted second-order N-electron valence state perturbation theory (SC-NEVPT2).
Acknowledgements
Y.R.P. and J.Y.-Z. were supported through the U.S. Department of Energy (DOE) under 2019030-SP DOE CalTech Sub S532207 (DE-SC0022089).